# A Factuality and Diversity Reconciled Decoding Method for Knowledge-Grounded Dialogue Generation

### v1

@@ Introduction
Knowledge-grounded dialogue generation aims to alleviate dull and unfaithful responses by infusing external knowledge into the input of dialogue models, and it encompasses two sub-tasks: knowledge selection and response generation. The hotspot of early research was primarily focused on how to enhance the performance of knowledge selection (Sun, Ren, and Ren 2023; Xu et al. 2022; Zhan et al. 2021; Yang et al. 2022; Meng et al. 2021; Kim, Ahn, and Kim 2020). With the substantial leap in the capabilities of generative models, the research focus gradually shifts to the response generation sub-task (Zhao et al. 2020a; Liu et al. 2021; Zhao et al. 2020b; Zheng, Milic-Frayling, and Zhou 2021). Ideally, an outstanding chat-bot should generate informative and truthful responses while maintaining the naturalistic phrasing and excellent interactivity (Dziri et al. 2022a).
@@
Hallucination pertains to the issue wherein the generated contents deviate from the user’s instruction, contradict the previously generated context, or contravene the established world (Zhang et al. 2023c). It presents severe risks to the practical applications in NLP. Existing works have conducted comprehensive explorations on hallucination from multiple viewpoints, including its origin (McKenna et al. 2023; Dziri et al. 2022b), detection (Zhang et al. 2023a; Manakul, Liusie, and Gales 2023; Fadeeva et al. 2023), and mitigation (Choi et al. 2023; Chuang et al. 2023; Li et al. 2023b). Numerous approaches and benchmarks regarding hallucination were proposed, such as question answering (Gao et al. 2023; Lin, Hilton, and Evans 2022), text summarization (Cao et al. 2020; Zhong et al. 2021), and dialogue generation (Li et al. 2023a; Wan et al. 2023; Chen et al. 2023; Sun et al. 2022). One effective solution to alleviate hallucination in QA is to rely on external knowledge as supplementary evidence, which provides more precise answers to user inquiries. Given the nature of open-domain conversations, a chit-chat agent should integrate external knowledge seamlessly into its responses, thereby attaining a balance between factuality and content diversity.
@@
Definition 3.1 $\mathcal{F}$): Given a response $y$, and external knowledge $\mathcal{K}=(k_{1},\ldots,k_{j})$ at turn $n$, we say that the response $y$ is faithful with respect to the external knowledge ($\mathcal{F}(\mathcal{K},y)$) if and only if the following condition holds:
@@
Definition 3.2 $\mathcal{F}^{*}$): Given a response $y$, we say that $y$ is factual ($\mathcal{F}^{*}(y)$) if and only if the following condition holds:
@@
Table 1: Automatic Evaluation results on the WoW dataset (LLaMA2-chat-7B). The best results are highlighted with bold. ”*” denotes that the improvement to the best baseline is statistically significant (t-test with $p$-value $<$ 0.01).
@@
Table 2: Automatic Evaluation results on the FaithDial and Holl-E dataset (LLaMA2-chat-7B). The best results are highlighted with bold. ”*” denotes that the improvement to the best baseline is statistically significant (t-test with $p$-value $<$ 0.01).
@@
$\mathbb{I}_{t}=\left\{\begin{array}[]{ll}1&\mbox{if}\ (\mathcal{F}^{c}_{t}>0)\;\vee\;(F^{k}_{t}-F^{c}_{t}<0),\\[4.30554pt] 0&\mbox{otherwise}.\\ \end{array}\right.$ (8)
@@
$\hat{p}_{c}(y_{t})=\left\{\begin{array}[]{ll}p_{c}(y_{t})/\hat{P}&\mbox{if}\ y_{t}\in\mathcal{V}^{(p)}\\[4.30554pt] 0&\mbox{otherwise},\\ \end{array}\right.$ (10)
@@ Appendix E Appendix E Details of the Prompts
Table 6: Automatic Evaluation results on the WoW dataset (LLaMA2-chat-13B). The best results are highlighted with bold. ”*” denotes that the improvement to the best baseline is statistically significant (t-test with $p$-value $<$ 0.01).
@@
Table 7: Automatic Evaluation results on the FaithDial and Holl-E dataset (LLaMA2-chat-13B). The best results are highlighted with bold. ”*” denotes that the improvement to the best baseline is statistically significant (t-test with $p$-value $<$ 0.01).
Figure 6: Human evaluation results on the Wizard of Wikipedia test seen set. The results are statistically significant with p-value $<$ 0.05, and Kappa ($\kappa$) falls between 0.4 and 0.6, suggesting moderate agreement among annotators.

### v2

@@ Introduction
Knowledge-grounded dialogue generation aims to alleviate dull and unfaithful responses by infusing external knowledge into the input of dialogue models, and it encompasses two sub-tasks: knowledge selection and response generation. The hotspot of early research was primarily focused on how to enhance the performance of knowledge selection (Sun, Ren, and Ren 2023; Xu et al. 2022; Zhan et al. 2021; Yang et al. 2022; Meng et al. 2021; Kim, Ahn, and Kim 2020; Wang et al. 2025; Yang et al. 2023b). With the substantial leap in the capabilities of generative models, the research focus gradually shifts to the response generation sub-task (Zhao et al. 2020a; Liu et al. 2021; Zhao et al. 2020b; Zheng, Milic-Frayling, and Zhou 2021; Chen et al. 2024). Ideally, an outstanding chat-bot should generate informative and truthful responses while maintaining the naturalistic phrasing and excellent interactivity (Dziri et al. 2022a).
@@
Hallucination pertains to the issue wherein the generated contents deviate from the user’s instruction, contradict the previously generated context, or contravene the established world (Zhang et al. 2023c). It presents severe risks to the practical applications in NLP. Existing works have conducted comprehensive explorations on hallucination from multiple viewpoints, including its origin (McKenna et al. 2023; Dziri et al. 2022b), detection (Zhang et al. 2023a; Manakul, Liusie, and Gales 2023; Fadeeva et al. 2023), and mitigation (Choi et al. 2023; Chuang et al. 2023; Li et al. 2023b; Yang et al. 2024). Numerous approaches and benchmarks regarding hallucination were proposed, such as question answering (Gao et al. 2023; Lin, Hilton, and Evans 2022), text summarization (Cao et al. 2020; Zhong et al. 2021), and dialogue generation (Li et al. 2023a; Wan et al. 2023; Chen et al. 2023; Sun et al. 2022; Wang et al. 2022). One effective solution to alleviate hallucination in QA is to rely on external knowledge as supplementary evidence, which provides more precise answers to user inquiries. Given the nature of open-domain conversations, a chit-chat agent should integrate external knowledge seamlessly into its responses, thereby attaining a balance between factuality and content diversity.
@@
Definition 3.1 Faithfulness($\mathcal{F}$): Given a response $y$, and external knowledge $\mathcal{K}=(k_{1},\ldots,k_{j})$ at turn $n$, we say that the response $y$ is faithful with respect to the external knowledge ($\mathcal{F}(\mathcal{K},y)$) if and only if the following condition holds:
@@
Definition 3.2 Factuality($\mathcal{F}^{*}$): Given a response $y$, we say that $y$ is factual ($\mathcal{F}^{*}(y)$) if and only if the following condition holds:
@@
Table 1: Automatic Evaluation results on the WoW dataset (LLaMA2-chat-7B). The best results are highlighted with bold. ”*” denotes that the improvement to the best baseline is statistically significant (t-test with $p$-value << 0.01).
@@
Table 2: Automatic Evaluation results on the FaithDial and Holl-E dataset (LLaMA2-chat-7B). The best results are highlighted with bold. ”*” denotes that the improvement to the best baseline is statistically significant (t-test with $p$-value << 0.01).
@@
[…]
@@
[…]
@@ Appendix E Appendix E Details of the Prompts
Table 6: Automatic Evaluation results on the WoW dataset (LLaMA2-chat-13B). The best results are highlighted with bold. ”*” denotes that the improvement to the best baseline is statistically significant (t-test with $p$-value << 0.01).
@@
Table 7: Automatic Evaluation results on the FaithDial and Holl-E dataset (LLaMA2-chat-13B). The best results are highlighted with bold. ”*” denotes that the improvement to the best baseline is statistically significant (t-test with $p$-value << 0.01).
Figure 6: Human evaluation results on the Wizard of Wikipedia test seen set. The results are statistically significant with p-value << 0.05, and Kappa ($\kappa$) falls between 0.4 and 0.6, suggesting moderate agreement among annotators.